OpenStudy (anonymous):

solving radical equations : solve for x -> root[x+10]=4 and root[3x+1]=12

OpenStudy (anonymous):

root[x+10]=4 /^2 x+10 = 16 x = 6 the second one: root[3x+1]=12 /^2 3x+1 = 144 3x= 143 x= 143/3